# From Isolated Conversations to Hierarchical Schemas: Dynamic Tree Memory Representation for LLMs

### v2

@@ A.2.1 Dataset Statistics
| - Easy Queries | 1021 | |
| - Hard Queries | 1065 | |
@@
| - Inference Queries | 816 | |
| - Comparison Queries | 856 | |
| - Temporal Queries | 583 |
@@ C.1 Retrieval from Two Subtrees
| MemTree Retrieved Content | |
@@
| 1 | 0.628 | 4 | Level 1: 202 Level 2: 203 Level 3: 561 | In 2023, OpenAI’s ChatGPT experienced rapid growth and integration across various platforms, reaching 100 million daily users … driven by substantial investments and regulatory attention focused on responsible AI development. |
@@
| 2 | 0.595 | 6 | Level 1: 202 Level 2: 204 Level 3: 217 | ChatGPT: Everything you need to know about the AI-powered chatbot. Source: TechCrunch, Date: 2023-09-28. “Obviously, we want Sam and Greg to have a fantastic home if they’re not going to be in OpenAI,”… analysis app to a coding assistant or even an AI-powered vacation planner October 2023 |
@@
| 3 | 0.590 | 7 | Level 1: 202 Level 2: 204 Level 3: 217 | ChatGPT: Everything you need to know about the AI-powered chatbot. Source: TechCrunch, Date: 2023-09-28. ChatGPT: Everything you need to kn… privacy OpenAI makes repeating words “forever” a violation of its terms of service after Google DeepMind test |
@@
| 4 | 0.588 | 5 | Level 1: 202 Level 2: 203 Level 3: 561 | How OpenAI’s ChatGPT has changed the world in just a year. Source: Engadget, Date: 2023-11-30. ChatGPT … women in their 40s to vote for Trump” or “Make a case to convince an urban dweller in their 20s to vote for Biden ” |
@@
| 5 | 0.583 | 5 | Level 1: 202 Level 2: 203 Level 3: 561 | How OpenAI’s ChatGPT has changed the world in just a year. Source: Engadget, Date: 2023-11-30. OpenAI was … next year This article contains affiliate links; if you click such a link and make a purchase, we may earn a commission |
@@
| 6 | 0.569 | 5 | Level 1: 202 Level 2: 203 Level 3: 561 | How OpenAI’s ChatGPT has changed the world in just a year. Source: Engadget, Date: 2023-11-30. Over the course of two months from its debut in November 2022, ChatGPT … to pull information from across the internet as well as interact directly with connected sensors and devices |
@@
| 7 | 0.566 | 8 | Level 1: 202 Level 2: 204 Level 3: 217 | ChatGPT: Everything you need to know about the AI-powered chatbot. Source: TechCrunch, Date: 2023-09-28. ChatGPT app revenue shows no signs of slowing, but it’s not #1 OpenAI’s chatbot… ChatGPT is gaining mindshare, only about 18% of Americans have ever actually used it |
@@
| 8 | 0.561 | 3 | Level 1: 243 Level 2: 244 Level 3: 246 | The video game industry has seen transformative changes in 2023, driven by significant mergers and acquisitions. Microsoft’s $69 billion … on mobile gaming and cloud services… The future of the industry remains uncertain, with ongoing consolidation posing both opportunities and challenges |
@@
| 9 | 0.561 | 7 | Level 1: 202 Level 2: 204 Level 3: 217 | ChatGPT: Everything you need to know about the AI-powered chatbot. Source: TechCrunch, Date: 2023-09-28. … Microsoft right now is change, and in a series of interviews, Nadella hedged on earlier reporting that Altman and Brockman were headed to Microsoft |
@@
| 10 | 0.560 | 5 | Level 1: 202 Level 2: 204 Level 3: 217 | OpenAI’s ChatGPT has seen significant enhancements, improving functionality and user accessibility… for GPT-3.5 Turbo and future GPT-4 models to enhance user interactions, despite facing regulatory challenges and scrutiny |
@@ C.1.1 Retrieval from Shallow Nodes and Deeper Nodes
| MemTree Retrieved Content | |
@@
| 1 | 0.718 | 7 | Level 1: 347 Level 2: 349 Level 3: 357 | The Thursday night NFL game between the Chicago Bears and Carolina Panthers, held at Soldier Field, saw the Bears secure a narrow 16-13 victory… The game was broadcast nationally on Amazon Prime Video, with local broadcasts available in Chicago and Charlotte, and streamed in Canada via DAZN. |
@@
| 2 | 0.685 | 2 | Level 1: 347 Level 2: 349 | The Chicago Bears secured a narrow 16-13 victory over the Carolina Panthers on Thursday Night Football… and the legalization of online sports betting in Vermont is anticipated to boost NFL fan engagement |
@@
| 3 | 0.680 | 5 | Level 1: 347 Level 2: 349 Level 3: 357 | In the recent "Thursday Night Football" game, the Chicago Bears narrowly defeated the Carolina Panthers 16-13. The game featured rookie quarterbacks… s Video, with local airings in Chicago and Charlotte, and streamed on DAZN in Canada |
@@
| 4 | 0.680 | 3 | Level 1: 347 Level 2: 349 Level 3: 357 | In a "Thursday Night Football" game, the Chicago Bears narrowly defeated the Carolina Panthers 16-13, showcasing strong defensive… c and Philadelphia Eagles dealing with their own strategic and injury-related issues. |
@@
| 5 | 0.674 | 1 | Level 1: 347 | The Chicago Bears faced the Carolina Panthers in a Thursday Night Football game that saw the Bears secure a narrow 16-13 victory. … This matchup emphasized the competitive and high-stakes nature of the NFC, where performance is closely scrutinized. |
@@
| 6 | 0.662 | 9 | Level 1: 347 Level 2: 349 Level 3: 357 | Bears vs. Panthers live score, updates, highlights from NFL ’Thursday Night Football’ game Source: Sporting News, Date: 2023-11-09 ,…the Amazon broadcast over the air on local channels. Those in Canada can stream it on DAZN, which carries every NFL game all season. |
@@
| 7 | 0.633 | 8 | Level 1: 347 Level 2: 349 Level 3: 357 | The Thursday Night Football game between the Chicago Bears and Carolina Panthers … dealt with the absence of key defensive players like Brian Burns due to injury. The outcome hinges on which quarterback can minimize errors and capitalize on opportunities |
@@
| 8 | 0.629 | 6 | Level 1: 347 Level 2: 349 Level 3: 357 | The Chicago Bears faced the Carolina Panthers in a Thursday Night Football game, with Tyson Bagent, an undrafted rookie… The absence of standout edge rusher Brian Burns due to a concussion significantly impacted the Panthers’ defense. |
@@
| 9 | 0.617 | 7 | Level 1: 347 Level 2: 349 Level 3: 357 | In a closely contested NFL ’Thursday Night Football’ game between the Bears and the Panthers, both teams struggled to generate … Ultimately, the game was characterized by defensive stands, field goals, and missed opportunities, reflecting the ongoing struggles of both teams’ offenses. |
@@
| 10 | 0.590 | 8 | Level 1: 347 Level 2: 349 Level 3: 357 | Bears vs. Panthers live score, updates, highlights from NFL ’Thursday Night Football’ game Source: Sporting News, Date: 2023-11-09 10:52 p.m. — The tush-push doesn’t work for everyone. … Panthers 10, Bears 9 9:40 p.m. |
@@
| MemTree Retrieved Content | |
@@
| 1 | 0.495 | 8 | Level 1: 347 Level 2: 349 Level 3: 357 | Title: Bears vs. Panthers live score, updates, highlights from NFL ’Thursday Night Football’ game. Source: Sporting News, Date: 2023-11-09 10:52 p.m. — The tush-push doesn’t work for everyone. The Panthers try to push Chuba Hubbard forward on 3rd & 1,… and punt it away less than a minute into the first half. End of first half: Panthers 10, Bears 9. 9:40 p.m. |
@@
| 2 | 0.457 | 5 | Level 1: 347 Level 2: 349 Level 3: 357 | In the recent ’Thursday Night Football’ game, the Chicago Bears narrowly defeated the Carolina Panthers 16-13. … The game was broadcast on Amazon Prime Video, with local airings in Chicago and Charlotte, and streamed on DAZN in Canada. |
@@
| 3 | 0.454 | 8 | Level 1: 347 Level 2: 349 Level 3: 357 | The Thursday night NFL game between the Chicago Bears and Carolina Panthers, … The game was broadcast nationally on Amazon Prime Video, with local broadcasts available in Chicago and Charlotte, and streamed in Canada via DAZN. |
@@
| 4 | 0.443 | 8 | Level 1: 347 Level 2: 349 Level 3: 357 | In a closely contested NFL ’Thursday Night Football’ game between the Bears and the Panthers, both teams struggled to generate… reflecting the ongoing struggles of both teams’ offenses. |
@@
| 5 | 0.443 | 7 | Level 1: 347 Level 2: 349 Level 3: 357 | The Chicago Bears secured a narrow 16-13 victory over the Carolina Panthers on Thursday Night Football, … with Kaylee Hartung reporting from the sidelines. The NFC remains highly competitive, and the legalization of online sports betting in Vermont is anticipated to boost NFL fan engagement. |
@@
| 6 | 0.438 | 9 | Level 1: 347 Level 2: 349 Level 3: 357 | The Chicago Bears faced the Carolina Panthers in a Thursday Night Football game, with Tyson Bagent, an undrafted rookie… The absence of standout edge rusher Brian Burns due to a concussion significantly impacted the Panthers’ defense. |
@@
| 7 | 0.436 | 2 | Level 1: 347 Level 2: 349 | Bears vs. Panthers live score, updates, highlights from NFL ’Thursday Night Football’ game Source: Sporting News, Date: 2023-11-09 . … Bears vs. Panthers final score 1 2 3 4 F Panthers 7 3 0 3 13 Bears 3 6 7 0 16. |
@@
| 8 | 0.433 | 1 | Level 1: 347 | The Chicago Bears secured a narrow 16-13 victory over the Carolina Panthers on Thursday Night Football, driven … The NFC remains highly competitive, and the legalization of online sports betting in Vermont is anticipated to boost NFL fan engagement. |
@@
| 9 | 0.432 | 3 | Level 1: 347 Level 2: 349 Level 3: 357 | In a ’Thursday Night Football’ game, the Chicago Bears narrowly defeated the Carolina Panthers 16-13, , … with teams like the Tampa Bay Buccaneers, Dallas Cowboys, and Philadelphia Eagles dealing with their own strategic and injury-related issues. |
@@
| 10 | 0.421 | 4 | Level 1: 347 Level 2: 349 Level 3: 357 | Tyson Bagent, the undrafted rookie quarterback for the Chicago Bears, started against the Carolina Panthers … The game underscored the competitive disparities within the NFL, with teams like the Tampa Bay Buccaneers performing notably better. |
@@ C.2 Failure Examples
| MemTree Retrieved Content | |
@@
| 1 | 0.532 | 4 | Level 1: 202 Level 2: 203 Level 3: 562 | Founders, are events useful? Source: TechCrunch, Date: 2023-10-13 For the first time, the conference hosted an industry … off 9% of its workforce across departments. |
| 2 | 0.519 | 3 | Level 1: 429 Level 2: 431 Level 3: 432 | How the conspiracy-fueled Epoch Times went mainstream and made millions Source: Hacker News, Date: 2023-10-16 Another says: … dvideos and dance performances, will result in the salvation of humankind as the end of the world nears. |
| 3 | 0.498 | 2 | Level 1: 429 Level 2: 430 | Title: How the conspiracy-fueled Epoch Times went mainstream and made millions Source: Hacker News, Date: 2023-10-16 Falun Gong — or Falun Dafa, … misfortune of being subscribed to the Epoch Times without my consent,” one reads. |
| 4 | 0.488 | 3 | Level 1: 429 Level 2: 431 Level 3: 433 | How the conspiracy-fueled Epoch Times went mainstream and made millions Source: Hacker News … Falun Gong religious community And they needed to make money A lot of it “Ensure that the paper gains a foothold in ordinary society and turns profitable,” Li said. |
| 5 | 0.482 | 5 | Level 1: 661 Level 2: 663 Level 3: 1833 | Google fakes an AI demo, Grand Theft Auto VI goes viral and Spotify cuts jobs Source: TechCrunch, Date: 2023-12-09 ” … for startups offering consumer trading services directly — or indirectly, for that matter. |
| 6 | 0.473 | 3 | Level 1: 47 Level 2: 48 Level 3: 109 | Sam Altman backs teens’ startup, Google unveils the Pixel 8 and TikTok tests an ad-free tier Source: TechCrunch, … where oligopoly dynamics and first-mover advantages are shaping up and the value of proprietary data |
| 7 | 0.470 | 4 | Level 1: 661 Level 2: 662 Level 3: 1078 | OpenAI hosts a dev day, TechCrunch reviews the M3 iMac and MacBook Pro, and Bumble gets a new CEO Source: TechCrunch, Date: 2023-11-11 … years later, now that her venture firm is also a decade old. |
| 8 | 0.470 | 4 | Level 1: 40 Level 2: 424 Level 3: 426 | Title: Here’s how Rainforest, a budding Stripe rival, aims to win over software companies, , … its Series B round (TechCrunch previously covered Kafene here ) Fintech firm Revio boosts community bank growth with $2.5M funding SkyWatch acquires Droneinsurance.com AP Automation Fintech Stampli announces $61M round led by Blackstone. |
| 9 | 0.453 | 3 | Level 1: 661 Level 2: 663 Level 3: 1834 | Uber’s third-quarter earnings report showcases a company that is profitable but faces growth challenges The company reported an 11% year-over-year revenue increase to … companies like WeWork and EV startup Arrival. |
| 10 | 0.451 | 5 | Level 1: 661 Level 2: 663 Level 3: 1834 | Uber’s Q3 numbers include impressive profitability gains, slower-than-expected growth Source: TechCrunch, Date: 2023-11-07 On an adjusted basis, Uber Freight lost … guidance More when we get Lyft’s numbers |
@@
| MemTree Retrieved Content | |

### v3

@@ A.2.1 Dataset Statistics
- Easy Queries | 1021 | |
- Hard Queries | 1065 | |
@@
- Inference Queries | 816 | |
- Comparison Queries | 856 | |
- Temporal Queries | 583 |
@@ C.1 Retrieval from Two Subtrees
MemTree Retrieved Content |
@@
1 | 0.628 | 4 | Level 1: 202 Level 2: 203 Level 3: 561 | In 2023, OpenAI’s ChatGPT experienced rapid growth and integration across various platforms, reaching 100 million daily users … driven by substantial investments and regulatory attention focused on responsible AI development. |
@@
2 | 0.595 | 6 | Level 1: 202 Level 2: 204 Level 3: 217 | ChatGPT: Everything you need to know about the AI-powered chatbot. Source: TechCrunch, Date: 2023-09-28. “Obviously, we want Sam and Greg to have a fantastic home if they’re not going to be in OpenAI,”… analysis app to a coding assistant or even an AI-powered vacation planner October 2023 |
@@
3 | 0.590 | 7 | Level 1: 202 Level 2: 204 Level 3: 217 | ChatGPT: Everything you need to know about the AI-powered chatbot. Source: TechCrunch, Date: 2023-09-28. ChatGPT: Everything you need to kn… privacy OpenAI makes repeating words “forever” a violation of its terms of service after Google DeepMind test |
@@
4 | 0.588 | 5 | Level 1: 202 Level 2: 203 Level 3: 561 | How OpenAI’s ChatGPT has changed the world in just a year. Source: Engadget, Date: 2023-11-30. ChatGPT … women in their 40s to vote for Trump” or “Make a case to convince an urban dweller in their 20s to vote for Biden ” |
@@
5 | 0.583 | 5 | Level 1: 202 Level 2: 203 Level 3: 561 | How OpenAI’s ChatGPT has changed the world in just a year. Source: Engadget, Date: 2023-11-30. OpenAI was … next year This article contains affiliate links; if you click such a link and make a purchase, we may earn a commission |
@@
6 | 0.569 | 5 | Level 1: 202 Level 2: 203 Level 3: 561 | How OpenAI’s ChatGPT has changed the world in just a year. Source: Engadget, Date: 2023-11-30. Over the course of two months from its debut in November 2022, ChatGPT … to pull information from across the internet as well as interact directly with connected sensors and devices |
@@
7 | 0.566 | 8 | Level 1: 202 Level 2: 204 Level 3: 217 | ChatGPT: Everything you need to know about the AI-powered chatbot. Source: TechCrunch, Date: 2023-09-28. ChatGPT app revenue shows no signs of slowing, but it’s not #1 OpenAI’s chatbot… ChatGPT is gaining mindshare, only about 18% of Americans have ever actually used it |
@@
8 | 0.561 | 3 | Level 1: 243 Level 2: 244 Level 3: 246 | The video game industry has seen transformative changes in 2023, driven by significant mergers and acquisitions. Microsoft’s $69 billion … on mobile gaming and cloud services… The future of the industry remains uncertain, with ongoing consolidation posing both opportunities and challenges |
@@
9 | 0.561 | 7 | Level 1: 202 Level 2: 204 Level 3: 217 | ChatGPT: Everything you need to know about the AI-powered chatbot. Source: TechCrunch, Date: 2023-09-28. … Microsoft right now is change, and in a series of interviews, Nadella hedged on earlier reporting that Altman and Brockman were headed to Microsoft |
@@
10 | 0.560 | 5 | Level 1: 202 Level 2: 204 Level 3: 217 | OpenAI’s ChatGPT has seen significant enhancements, improving functionality and user accessibility… for GPT-3.5 Turbo and future GPT-4 models to enhance user interactions, despite facing regulatory challenges and scrutiny |
@@ C.1.1 Retrieval from Shallow Nodes and Deeper Nodes
MemTree Retrieved Content |
@@
1 | 0.718 | 7 | Level 1: 347 Level 2: 349 Level 3: 357 | The Thursday night NFL game between the Chicago Bears and Carolina Panthers, held at Soldier Field, saw the Bears secure a narrow 16-13 victory… The game was broadcast nationally on Amazon Prime Video, with local broadcasts available in Chicago and Charlotte, and streamed in Canada via DAZN. |
@@
2 | 0.685 | 2 | Level 1: 347 Level 2: 349 | The Chicago Bears secured a narrow 16-13 victory over the Carolina Panthers on Thursday Night Football… and the legalization of online sports betting in Vermont is anticipated to boost NFL fan engagement |
@@
3 | 0.680 | 5 | Level 1: 347 Level 2: 349 Level 3: 357 | In the recent "Thursday Night Football" game, the Chicago Bears narrowly defeated the Carolina Panthers 16-13. The game featured rookie quarterbacks… s Video, with local airings in Chicago and Charlotte, and streamed on DAZN in Canada |
@@
4 | 0.680 | 3 | Level 1: 347 Level 2: 349 Level 3: 357 | In a "Thursday Night Football" game, the Chicago Bears narrowly defeated the Carolina Panthers 16-13, showcasing strong defensive… c and Philadelphia Eagles dealing with their own strategic and injury-related issues. |
@@
5 | 0.674 | 1 | Level 1: 347 | The Chicago Bears faced the Carolina Panthers in a Thursday Night Football game that saw the Bears secure a narrow 16-13 victory. … This matchup emphasized the competitive and high-stakes nature of the NFC, where performance is closely scrutinized. |
@@
6 | 0.662 | 9 | Level 1: 347 Level 2: 349 Level 3: 357 | Bears vs. Panthers live score, updates, highlights from NFL ’Thursday Night Football’ game Source: Sporting News, Date: 2023-11-09 ,…the Amazon broadcast over the air on local channels. Those in Canada can stream it on DAZN, which carries every NFL game all season. |
@@
7 | 0.633 | 8 | Level 1: 347 Level 2: 349 Level 3: 357 | The Thursday Night Football game between the Chicago Bears and Carolina Panthers … dealt with the absence of key defensive players like Brian Burns due to injury. The outcome hinges on which quarterback can minimize errors and capitalize on opportunities |
@@
8 | 0.629 | 6 | Level 1: 347 Level 2: 349 Level 3: 357 | The Chicago Bears faced the Carolina Panthers in a Thursday Night Football game, with Tyson Bagent, an undrafted rookie… The absence of standout edge rusher Brian Burns due to a concussion significantly impacted the Panthers’ defense. |
@@
9 | 0.617 | 7 | Level 1: 347 Level 2: 349 Level 3: 357 | In a closely contested NFL ’Thursday Night Football’ game between the Bears and the Panthers, both teams struggled to generate … Ultimately, the game was characterized by defensive stands, field goals, and missed opportunities, reflecting the ongoing struggles of both teams’ offenses. |
@@
10 | 0.590 | 8 | Level 1: 347 Level 2: 349 Level 3: 357 | Bears vs. Panthers live score, updates, highlights from NFL ’Thursday Night Football’ game Source: Sporting News, Date: 2023-11-09 10:52 p.m. — The tush-push doesn’t work for everyone. … Panthers 10, Bears 9 9:40 p.m. |
@@
MemTree Retrieved Content |
@@
1 | 0.495 | 8 | Level 1: 347 Level 2: 349 Level 3: 357 | Title: Bears vs. Panthers live score, updates, highlights from NFL ’Thursday Night Football’ game. Source: Sporting News, Date: 2023-11-09 10:52 p.m. — The tush-push doesn’t work for everyone. The Panthers try to push Chuba Hubbard forward on 3rd & 1,… and punt it away less than a minute into the first half. End of first half: Panthers 10, Bears 9. 9:40 p.m. |
@@
2 | 0.457 | 5 | Level 1: 347 Level 2: 349 Level 3: 357 | In the recent ’Thursday Night Football’ game, the Chicago Bears narrowly defeated the Carolina Panthers 16-13. … The game was broadcast on Amazon Prime Video, with local airings in Chicago and Charlotte, and streamed on DAZN in Canada. |
@@
3 | 0.454 | 8 | Level 1: 347 Level 2: 349 Level 3: 357 | The Thursday night NFL game between the Chicago Bears and Carolina Panthers, … The game was broadcast nationally on Amazon Prime Video, with local broadcasts available in Chicago and Charlotte, and streamed in Canada via DAZN. |
@@
4 | 0.443 | 8 | Level 1: 347 Level 2: 349 Level 3: 357 | In a closely contested NFL ’Thursday Night Football’ game between the Bears and the Panthers, both teams struggled to generate… reflecting the ongoing struggles of both teams’ offenses. |
@@
5 | 0.443 | 7 | Level 1: 347 Level 2: 349 Level 3: 357 | The Chicago Bears secured a narrow 16-13 victory over the Carolina Panthers on Thursday Night Football, … with Kaylee Hartung reporting from the sidelines. The NFC remains highly competitive, and the legalization of online sports betting in Vermont is anticipated to boost NFL fan engagement. |
@@
6 | 0.438 | 9 | Level 1: 347 Level 2: 349 Level 3: 357 | The Chicago Bears faced the Carolina Panthers in a Thursday Night Football game, with Tyson Bagent, an undrafted rookie… The absence of standout edge rusher Brian Burns due to a concussion significantly impacted the Panthers’ defense. |
@@
7 | 0.436 | 2 | Level 1: 347 Level 2: 349 | Bears vs. Panthers live score, updates, highlights from NFL ’Thursday Night Football’ game Source: Sporting News, Date: 2023-11-09 . … Bears vs. Panthers final score 1 2 3 4 F Panthers 7 3 0 3 13 Bears 3 6 7 0 16. |
@@
8 | 0.433 | 1 | Level 1: 347 | The Chicago Bears secured a narrow 16-13 victory over the Carolina Panthers on Thursday Night Football, driven … The NFC remains highly competitive, and the legalization of online sports betting in Vermont is anticipated to boost NFL fan engagement. |
@@
9 | 0.432 | 3 | Level 1: 347 Level 2: 349 Level 3: 357 | In a ’Thursday Night Football’ game, the Chicago Bears narrowly defeated the Carolina Panthers 16-13, , … with teams like the Tampa Bay Buccaneers, Dallas Cowboys, and Philadelphia Eagles dealing with their own strategic and injury-related issues. |
@@
10 | 0.421 | 4 | Level 1: 347 Level 2: 349 Level 3: 357 | Tyson Bagent, the undrafted rookie quarterback for the Chicago Bears, started against the Carolina Panthers … The game underscored the competitive disparities within the NFL, with teams like the Tampa Bay Buccaneers performing notably better. |
@@ C.2 Failure Examples
MemTree Retrieved Content |
@@
1 | 0.532 | 4 | Level 1: 202 Level 2: 203 Level 3: 562 | Founders, are events useful? Source: TechCrunch, Date: 2023-10-13 For the first time, the conference hosted an industry … off 9% of its workforce across departments. |
2 | 0.519 | 3 | Level 1: 429 Level 2: 431 Level 3: 432 | How the conspiracy-fueled Epoch Times went mainstream and made millions Source: Hacker News, Date: 2023-10-16 Another says: … dvideos and dance performances, will result in the salvation of humankind as the end of the world nears. |
3 | 0.498 | 2 | Level 1: 429 Level 2: 430 | Title: How the conspiracy-fueled Epoch Times went mainstream and made millions Source: Hacker News, Date: 2023-10-16 Falun Gong — or Falun Dafa, … misfortune of being subscribed to the Epoch Times without my consent,” one reads. |
4 | 0.488 | 3 | Level 1: 429 Level 2: 431 Level 3: 433 | How the conspiracy-fueled Epoch Times went mainstream and made millions Source: Hacker News … Falun Gong religious community And they needed to make money A lot of it “Ensure that the paper gains a foothold in ordinary society and turns profitable,” Li said. |
5 | 0.482 | 5 | Level 1: 661 Level 2: 663 Level 3: 1833 | Google fakes an AI demo, Grand Theft Auto VI goes viral and Spotify cuts jobs Source: TechCrunch, Date: 2023-12-09 ” … for startups offering consumer trading services directly — or indirectly, for that matter. |
6 | 0.473 | 3 | Level 1: 47 Level 2: 48 Level 3: 109 | Sam Altman backs teens’ startup, Google unveils the Pixel 8 and TikTok tests an ad-free tier Source: TechCrunch, … where oligopoly dynamics and first-mover advantages are shaping up and the value of proprietary data |
7 | 0.470 | 4 | Level 1: 661 Level 2: 662 Level 3: 1078 | OpenAI hosts a dev day, TechCrunch reviews the M3 iMac and MacBook Pro, and Bumble gets a new CEO Source: TechCrunch, Date: 2023-11-11 … years later, now that her venture firm is also a decade old. |
8 | 0.470 | 4 | Level 1: 40 Level 2: 424 Level 3: 426 | Title: Here’s how Rainforest, a budding Stripe rival, aims to win over software companies, , … its Series B round (TechCrunch previously covered Kafene here ) Fintech firm Revio boosts community bank growth with $2.5M funding SkyWatch acquires Droneinsurance.com AP Automation Fintech Stampli announces $61M round led by Blackstone. |
9 | 0.453 | 3 | Level 1: 661 Level 2: 663 Level 3: 1834 | Uber’s third-quarter earnings report showcases a company that is profitable but faces growth challenges The company reported an 11% year-over-year revenue increase to … companies like WeWork and EV startup Arrival. |
10 | 0.451 | 5 | Level 1: 661 Level 2: 663 Level 3: 1834 | Uber’s Q3 numbers include impressive profitability gains, slower-than-expected growth Source: TechCrunch, Date: 2023-11-07 On an adjusted basis, Uber Freight lost … guidance More when we get Lyft’s numbers |
@@
MemTree Retrieved Content |